OpenStudy (anonymous):

how do you evaluate the integral of xe^(-x^2)dx on the integral (0,infinity)?

OpenStudy (anonymous):

sub u=-x^2 so du=-2xdx

OpenStudy (anonymous):

1/2\[Limit as R->\infty of 1/2 \int\limits_{0}^{R} \]

OpenStudy (anonymous):

ah.. lol

OpenStudy (anonymous):

I would do sub: u=x^2/2, so du=u and integral ( e^u du)... should be easy from this point

OpenStudy (anonymous):

are you good with the improper integral part of it?

OpenStudy (anonymous):

Let me know what your answer is

OpenStudy (anonymous):

give me a min

OpenStudy (anonymous):

Okay. Medal plz

OpenStudy (anonymous):

Im getting: integral (e^(-u)du)=-e(-u) now: e^0=1 lim e(-u) when u is going to infinity = lim (1/ e^u) =0 my answer would be 1